Volume 6, 1994 – Volume 10, 1998

Title Index

Are There Field Dependent Formulas for the Roots to Polynomials?, Robert Cacioppo, Vol. 7, 10.

Applications of Colorability to Links, Wild Knots and Non–Compact Knots, Ollie Nanyes, Vol. 6, 124.

Base Negative Ten, Andrea Roth-

bart, Vol. 6, 91. Bounded and L^{p}–Solutions to a

Generalized Lienard Equation with Integrable Forcing Term, Allan Kroopnick, Vol. 10, 15.

Changepoint Detection Using Nonparametric Procedures, Sivanandan Balakumar, Vol. 9, 178.

A Characterization of Ultrafilters in Complementary Topologies, Rahim G. Karimpour, Vol. 9, 95.

The Circumcircle and the Line at Infinity, Clark Kimberling and Peter Yff, Vol. 9, 3.

Complex Continued Fractions: An Undergraduate Research Problem Proposal, Timothy P. Keller, Vol. 10, 147.

Composite Rules for Improper Integrals, Robert L. Doucette, Vol. 10, 83.

A Composition Problem Involving Analytic Functions, Paul E. Fishback, Vol. 10, 95.

Concurrency Theorems for Penta-

grams, Larry Hoehn, Vol. 8, 125. Constructing Pythagorean Triple

Preserving Matrices, Leonard Palmer, Mangho Ahuja, and Mohan Tikoo, Vol. 10, 159.

Cubic and Quartic Residues Modulo A Prime, Joseph B. Dence and Thomas P. Dence, Vol. 7, 24.

Some Criteria for the Homotopy Method for the Tridiagonal Eigenproblem, Kuiyuan Li, Vol 6, 154.

On the Decomposition of the Unit Interval, Vladimir Drobot, Vol. 8, 11.

On Dense Metrizable Subspaces of Topological Spaces, Ollie Nanyes, Vol. 7, 123.

A Development of Euler Numbers,

Joseph B. Dence, Vol. 9, 148. Digital Soup, Richard L. Francis,

Vol. 7, 57. On a Diophantine Equation Relat-

ed to Symmetric Design, H. W. Lu, Harold Bowman, and Peter J.-S. Shiue, Vol. 9, 133.

On Diophantine Equations x^{2} − dy^{2} = AB, Yungchen Cheng, Vol. 7, 68.

Entire Functions of Order One and Infinite Type, Badih Ghusayni, Vol. 10, 20.

Evaluating a Family of Integrals,

Russell Euler, Vol. 9, 156. On the Evaluation of Multiple In-

tegrals, K. L. D. Gunawardena, Vol. 6, 29.

Finding Pythagorean Triple Preserving Matrices, Leonard Palmer, Mangho Ahuja, and Mohan Tikoo, Vol. 10, 99.

Finite Presentations of Subgroups of Graph Groups, Joshua Levy, Cameron Parker, and Leonard VanWyk, Vol. 10, 70.

From the Legacy of Pythagoras, Herta T. Freitag, Vol. 8, 55.

George Green, Applied Mathemati-

cian, Mark Sand, Vol. 8, 22. Half–Life, Bone Lengths, Euler’s

Formula, and Constructivism: Ways to Integrate Mathematics and Science, Marvin Harrell and Kenneth Thompson, Vol. 9, 160.

On Hermitian–Invertors, Richard D. Hill and Joseph R. Siler, Vol. 8, 14.

Herta T. Freitag – Profile, Vol. 8,

54. An Historical Contradiction, David

L. Clements, Vol. 8, 82. Homotopy Method for the Sin-

gular Symmetric Tridiagonal Eigenproblem, Kuiyuan Li and Tien-Yien Li, Vol. 6, 34.

Homotopy Understanding of Iterative Methods, Noah H. Rhee, Vol. 6, 78.

The Ideal Structure of Z*Z, Tilak

de Alwis, Vol. 6, 116. Independent Random Variables on

the Unit Interval, James Foran and Lee Hart, Vol. 6, 144.

Isosceles Triangles Equal in Perimeter and Area, Paul Yiu, Vol. 10, 106.

Iterations on Convex Quadrilaterals, Tamela Underwood and Mangho Ahuja, Vol. 6, 10.

A Linear Programming Transformation of an Inconsistent System of Linear Equations into a Consistent System, Mohammad Fatehi, Miguel Paredes, Richard Hinthorn, and Hushang Poorkarimi, Vol. 7, 3.

Links of Certain Singularities,

Mathew Timm, Vol. 7, 76. Mixed Insurance Risk Models, Liansheng Chen and Jinhua Tao, Vol. 8, 3.

Monoids Connected with Euler’s Diophantine Equation, Aleksander Grytczuk, Vol. 10, 91. A Necessary and Sufficient Condition for Twin Primes, Joseph B. Dence and Thomas P. Dence, Vol. 7, 129.

A New Generalization of Reed–Muller Codes, Todd D. Vance, Vol. 9, 79.

Note on Birational Extensions in D–Dimension, Mee–Kyoung Kim, Vol. 10, 153.

A Note on Linear Differential Equations with Constant Coefficients, Richard E. Bayne, James E. Joseph, Myung H. Kwack, and Thomas H. Lawson, Vol. 10, 41. A Note on Module Homomorphism-

s, John Koker, Vol. 6, 69. A Note on a Pythagorean Ring,

Aleksander Grytczuk, Vol. 9, 83. A Note on the Sum ∑ 1/w_{k2^{n}}, Stanley Rabinowitz, Vol. 10, 141. An Operator Inequality, Derming

Wang, Vol. 7, 17. Pairwise Locally Compact Biconvergence Spaces, Jamuna P. Ambasht, Vol. 9, 28.

Parabolas in Taxicab Geometry, Y. Phoebe Ho and Yan Liu, Vol. 8, 63.

A Perplexing Finite Continued Fraction, Stanley Rabinowitz, Vol. 6, 2.

A Polynomial Curve of Constant Width, Stanley Rabinowitz, Vol. 9, 23.

Presentations of Subgroups of Artin Groups, Jennifer Becker, Matthew Horak, and Leonard VanWyk, Vol. 10, 3.

Probable Prime Predicaments,

Richard L. Francis, Vol. 7, 132. Products and Sums With Applica-

tions, Badih Ghusayni, Vol. 9, 90.

The Rational Zero Theorem Extended, Leonard L. Palmer, Vol. 7, 20.

A Relationship Between Chords and the Radius of a Circle, Russell Euler, Vol. 9, 102.

A Relationship Between Quadratic Equations, Right Triangles and Fibonacci Type Sequences, Leonard L. Palmer, Vol. 7, 119.

A Renaissance of Geometric Constructions, Richard L. Francis, Vol. 8, 113.

Residues – Part II. Congruences Modulo Powers of 2, Joseph B. Dence and Thomas P. Dence, Vol. 8, 26.

Residues – Part III. Congruences to General Composite Moduli, Joseph B. Dence and Thomas P. Dence, Vol. 9, 72

The Riemann-Stieltjes Integral, R.

E. Bradley, Vol. 6, 20. A Ring of Pythagorean Triples,

Bryan Dawson, Vol. 6, 72. Sequences of Mappings and the Ex-

pansive Property, Fred Worth, Vol. 8, 73.

Sequential G_{δ}–Sets and Measures,

Roger H. Marty, Vol. 7, 54. Set Valued Mappings on the Reals,

Kandasamy Muthuvel, Vol. 6, 139.

Sixes and Sevens, Carl Pomerance, Vol. 6, 62.

Some Characterizations of Perfect Numbers, Joe Flowers, Vol. 7, 104.

Some Improved Results on B–Rings, Amir M. Rahimi, Vol. 9, 167.

Some Representations of ζ(3), Badih Ghusayni, Vol. 10, 169. Stable Range in Formal Power Series with any Number of Indeterminates, Amir M. Rahimi, Vol. 10, 28.

Stable Rings and Side Divisors, Amir M. Rahimi, Vol. 10, 112. A Student’s Probability of Finding a Basis, Bryan Dawson, Vol. 10, 37.

A Topological Proof of the Cayley–Hamilton Theorem, Jeffrey A. Rosoff, Vol. 7, 63. Two Solutions of One Problem of B. S. Popov, Kostadin G. Trenc̆evski and Ice B. Risteski, Vol. 9, 140.

Subaddition, David Choate, Vol. 9,

170. Ultrafilters and Topological En-

tropy of Complementary Topologies, Rahim G. Karimpour, Vol. 7, 32.

On Weak* Separable Subsets of Dual Banach Spaces, Joe Howard, Vol. 7, 116.

Zero–Dimensional Nearness Spaces and Extensions of Topological Spaces, Rhonda L. McKee, Vol. 6, 64.

Author Index Ahuja, Mangho see Underwood,

Tamela, Vol. 6, 10. Ahuja, Mangho see Palmer,

Leonard L., Vol. 10, 99. Ahuja, Mangho see Palmer,

Leonard, Vol. 10, 159. Ambasht, Jamuna P., Pairwise Lo-

cally Compact Biconvergence Spaces, Vol. 9, 28.

Balakumar, Sivanandan, Changepoint Detection Using Nonparametric Procedures, Vol. 9, 178.

Bayne, Richard E., James E.

Joseph, Myung H. Kwack, and Thomas H. Lawson, A Note on Linear Differential Equations with Constant Coefficients, Vol. 10, 41.

Becker, Jennifer, Matthew Horak, and Leonard VanWyk, Presentations of Subgroups of Artin Groups, Vol. 10, 3.

Bowman, Harold see Lu, H. W.,

Vol. 9, 133. Bradley, R. E., The

Riemann-Stieltjes Integral, Vol. 6, 20.

Cacioppo, Robert, Are There Field Dependent Formulas for the Roots to Polynomials?, Vol. 7, 10.

Chen, Liansheng and Jinhua Tao, Mixed Insurance Risk Models, Vol. 8, 3.

Cheng, Yungchen, On Diophantine Equations x^{2} − dy^{2} = AB, Vol. 7, 68.

Choate, David, Subaddition, Vol.

9, 170. Clements, David L., An Historical

Contradiction, Vol. 8, 82.

Dawson, Bryan, A Ring of

Pythagorean Triples, Vol. 6, 72. Dawson, Bryan, A Student’s Probability of Finding a Basis, Vol. 10, 37.

de Alwis, Tilak, The Ideal Structure of Z*Z, Vol. 6, 116. Dence, Joseph B., A Development of Euler Numbers, Vol. 9, 148. Dence, Joseph B. and Thomas P. Dence, Cubic and Quartic Residues Modulo A Prime, Vol. 7, 24.

Dence, Joseph B. and Thomas P. Dence, A Necessary and Sufficient Condition for Twin Primes, Vol. 7, 129.

Dence, Joseph B. and Thomas P. Dence, Residues – Part II. Congruences Modulo Powers of 2, Vol. 8, 26.

Dence, Joseph B. and Thomas P. Dence, Residues – Part III. Congruences to General Composite Moduli, Vol. 9, 72.

Dence, Thomas P. see Dence, Joseph B., Vol. 7, 24. Dence, Thomas P. see Dence, Joseph B., Vol. 7, 129. Dence, Thomas P. see Dence, Joseph B., Vol. 8, 26. Dence, Thomas P. see Dence, Joseph B., Vol. 9, 72.

Doucette, Robert L., Composite Rules for Improper Integrals, Vol. 10, 83.

Drobot, Vladimir, On the Decomposition of the Unit Interval, Vol. 8, 11.

Euler, Russell, Evaluating a Family of Integrals, Vol. 9, 156.

Euler, Russell, A Relationship Between Chords and the Radius of a Circle, Vol. 9, 102.

Fatehi, Mohammad, Miguel Parades, Richard Hinthorn, and Hushang Poorkarimi, A Linear Programming Transformation of an Inconsistent System of Linear Equations into a Consistent System, Vol 7, 3.

Fishback, Paul E., A Composition Problem Involving Analytic Functions, Vol. 10, 95.

Flowers, Joe, Some Characterizations of Perfect Numbers, Vol. 7, 104.

Foran, James and Lee Hart, Independent Random Variables on the Unit Interval, Vol. 6, 144.

Francis, Richard L., Digital Soup,

Vol. 7, 57. Francis, Richard L., Probable

Prime Predicaments, Vol. 7, 132. Francis, Richard L., A Renaissance

of Geometric Constructions, Vol. 8, 113.

Freitag, Herta T., From the Legacy

of Pythagoras, Vol. 8, 55. Freitag, Herta T., Herta T. Freitag

– Profile, Vol. 8, 54. Ghusayni, Badih, Products and

Sums With Applications, Vol. 9, 90.

Ghusayni, Badih, Entire Functions of Order One and Infinite Type, Vol. 10, 20.

Ghusayni, Badih, Some Representa-

tions of ζ(3), Vol. 10, 169. Grytczuk, Aleksander, Note on a

Pythagorean Ring, Vol. 9, 83. Grytczuk, Aleksander, Monoids

Connected with Euler’s Diophan-

tine Equation, Vol. 10, 91. Gunawardena, K. L. D., On the

Evaluation of Multiple Integrals, Vol. 6, 29.

Harrell, Marvin and Kenneth Thompson, Half–Life, Bone Lengths, Euler’s Formula, and Constructivism: Ways to Integrate Mathematics and Science, Vol. 9, 160.

Hart, Lee see Foran, James, Vol. 6,

144. Hill, Richard D. and Joseph R.

Siler, On Hermitian–Invertors, Vol. 8, 14.

Ho, Y. Phoebe and Yan Liu, Parabolas in Taxicab Geometry, Vol. 8, 63.

Hoehn, Larry, Concurrency Theo-

rems for Pentagrams, Vol. 8, 125. Horak, Matthew see Becker, Jen-

nifer, Vol. 10, 3. Howard, Joe, On Weak* Separable

Subsets of Dual Banach Spaces, Vol. 7, 116.

Hinthorn, Richard see Fatehi, Mo-

hammad, Vol. 7, 3. Joseph, James E. see Bayne,

Richard E., Vol. 10, 41. Karimpour, Rahim G., A Charac-

terization of Ultrafilters in Complementary Topologies, Vol. 9, 95.

Karimpour, Rahim G., Ultrafilters and Topological Entropy of Complementary Topologies, Vol. 7, 32.

Keller, Timothy P., Complex Continued Fractions: An Undergraduate Research Problem Proposal, Vol. 10, 147.

Kim, Mee–Kyoung, Note on Birational Extensions in D–Dimension, Vol. 10, 153.

Kimberling, Clark and Peter Yff, The Circumcircle and the Line at Infinity, Vol. 9, 3.

Koker, John, A Note on Module

Homomorphisms, Vol. 6, 69. Kroopnick, Allan, Bounded and

L^{p}–Solutions to a Generalized Lienard Equation with Integrable Forcing Term, Vol. 10, 15.

Kwack, Myung H. see Bayne,

Richard E., Vol. 10, 41. Lawson, Thomas H. see Bayne,

Richard E., Vol. 10, 41. Levy, Joshua, Cameron Parker, and

Leonard VanWyk, Finite Presentations of Subgroups of Graph Groups, Vol. 10, 70.

Li, Kuiyuan and Tien-Yien Li, Homotopy Method for the Singular Symmetric Tridiagonal Eigenproblem, Vol. 6, 34.

Li, Kuiyuan, Some Criteria for the Homotopy Method for the Tridiagonal Eigenproblem, Vol. 6, 154.

Li, Tien-Yien see Li, Kuiyuan, Vol.

6, 34. Liu, Yan see Ho, Y. Phoebe, Vol.

8, 63. Lu, H. W., Harold Bowman and

Peter J.-S. Shiue, On a Diophantine Equation Related to Symmetric Design, Vol. 9, 133.

Marty, Roger H., Sequential G_{δ}–Sets and Measures, Vol. 7, 54.

McKee, Rhonda L., Zero–Dimensional Nearness Spaces and Extensions of Topological Spaces, Vol. 6, 64.

Muthuvel, Kandasamy, Set Valued Mappings on the Reals, Vol. 6, 139.

Nanyes, Ollie, Applications of Colorability to Links, Wild Knots and Non–Compact Knots, Vol. 6, 124.

Nanyes, Ollie, On Dense Metrizable Subspaces of Topological Spaces, Vol. 7, 123.

Palmer, Leonard L., The Rational Zero Theorem Extended, Vol. 7, 20.

Palmer, Leonard L., A Relationship Between Quadratic Equations, Right Triangles and Fibonacci Type Sequences, Vol. 7, 119.

Palmer, Leonard L., Mangho Ahuja, and Mohan Tikoo, Finding Pythagorean Triple Preserving Matrices, Vol. 10, 99.

Palmer, Leonard, Mangho Ahuja, and Mohan Tikoo, Constructing Pythagorean Triple Preserving Matrices, Vol. 10, 159.

Parades, Miguel see Fatehi, Mo-

hammad, Vol. 7, 3. Parker, Cameron see Levy, Joshua,

Vol. 10, 70. Pomerance, Carl, Sixes and Sevens,

Vol. 6, 62. Poorkarimi, Hushang see Fatehi,

Mohammad, Vol. 7, 3. Rabinowitz, Stanley, A Perplexing

Finite Continued Fraction, Vol. 6, 2.

Rabinowitz, Stanley, A Polynomial Curve of Constant Width, Vol. 9, 23.

Rabinowitz, Stanley, A Note on the

Sum ∑ 1/w_{k2^{n}}, Vol. 10, 141. Rahimi, Amir M., Some Improved

Results on B–Rings, Vol. 9, 167. Rahimi, Amir M., Stable Range in

Formal Power Series with any Number of Indeterminates, Vol. 10, 28.

Rahimi, Amir M., Stable Rings and

Side Divisors, Vol. 10, 112. Rhee, Noah H., Homotopy Under-

standing of Iterative Methods, Vol. 6, 78.

Risteski, Ice B. see Trenc̆evski,

Kostadin G., Vol. 9, 140. Rosoff, Jeffrey A., A Topological

Proof of the Cayley–Hamilton Theorem, Vol. 7, 63.

Rothbart, Andrea, Base Negative

Ten, Vol. 6, 91. Sand, Mark, George Green, Applied

Mathematician, Vol. 8, 22. Shiue, Peter J.-S. see Lu, H. W.,

Vol. 9, 133. Siler, Joseph R. see Hill, Richard

D., Vol. 8, 14. Tao, Jinhua see Chen, Liansheng,

Vol. 8, 3. Thompson, Kenneth see Harrell,

Marvin, Vol. 9, 160. Tikoo, Mohan see Palmer, Leonard

L., Vol. 10, 99. Tikoo, Mohan see Palmer, Leonard,

Vol. 10, 159. Timm, Mathew, Links of Certain

Singularities, Vol. 7, 76. Trenc̆evski, Kostadin G. and Ice B.

Risteski, Two Solutions of One Problem of B. S. Popov, Vol. 9, 140.

Underwood, Tamela and Mangho Ahuja, Iterations on Convex

Quadrilaterals, Vol. 6, 10. Vance, Todd D., A New Generalization of Reed–Muller Codes, Vol. 9, 79.

VanWyk, Leonard see Becker, Jen-

nifer, Vol. 10, 3. VanWyk, Leonard see Levy, Joshua, Vol. 10, 70.

Wang, Derming, An Operator In-

equality, Vol. 7, 17. Worth, Fred, Sequences of Mappings and the Expansive Property, Vol. 8, 73.

Yff, Peter see Kimberling, Clark,

Vol. 9, 3. Yiu, Paul, Isosceles Triangles Equal in Perimeter and Area, Vol. 10, 106.

Problems Proposed Azarian, Mohammad K., Vol. 6,

98. Azarian, Mohammad K., Vol. 9,

34. Azarian, Mohammad K., Vol. 9,

105. Azarian, Mohammad K., Vol. 9,

184. Azarian, Mohammad K., Vol. 10,

176. Beltramo, Alvin, Vol. 6, 47. Chen, Xuming and Kenneth Daven-

port, Vol. 10, 118. Cooper, Curtis and Robert E.

Kennedy, Vol. 6, 160. Cooper, Curtis and Robert E.

Kennedy, Vol. 7, 40. Cooper, Curtis and Robert E.

Kennedy, Vol. 7, 87. Cooper, Curtis see Kennedy,

Robert E., Vol. 7, 139. Cooper, Curtis and Robert E.

Kennedy, Vol. 7, 140. Cooper, Curtis and Robert E.

Kennedy, Vol. 8, 90. Cooper, Curtis see Rabinowitz, Vol.

8, 90. Davenport, Kenneth, Vol. 9, 35. Davenport, Kenneth, Vol. 9, 105. Davenport, Kenneth, Vol. 9, 184. Davenport, Kenneth, Vol. 10, 46. Davenport, Kenneth, Vol. 10, 47. Davenport, Kenneth, Vol. 10, 117. Davenport, Kenneth see Chen, X-

uming, Vol. 10, 118. Dawson, Bryan, Vol. 8, 136. Dence, Joseph B., Vol. 6, 48. Dence, Joseph B., Vol. 8, 36. Dence, Joseph B., Vol. 8, 37. Dence, Joseph B., Vol. 8, 89. Dence, Joseph B., Vol. 9, 106.

Dence, Joseph B. see Dence,

Thomas P., Vol. 9. Dence, Thomas P. and Joseph B.

Dence, Vol. 9, 35. Etheridge, James see Wheatley,

Vol. 7. Euler, Russell and Jawad Sadek,

Vol. 10, 47. Euler, Russell and Jawad Sadek,

Vol. 10, 118. Freitag, Herta T., Vol. 6, 98. Freitag, Herta T., Vol. 6, 160. Freitag, Herta T., Vol. 6, 160. Freitag, Herta T., Vol. 7, 39. Freitag, Herta T., Vol. 7, 40. Freitag, Herta T., Vol. 8, 37. Freitag, Herta T., Vol. 8, 91. Freitag, Herta T., Vol. 8, 134. Freitag, Herta T., Vol. 8, 135. Freitag, Herta T., Vol. 8, 135. Freitag, Herta T., Vol. 9, 185. Gupton, Ronnie, Larry Hoehn, and

Jim Ridenhour, Vol. 6, 98. Hoehn, Larry, Vol. 7, 40. Hoehn, Larry, Vol. 10, 118. Hoehn, Larry, Vol. 10, 177. Hoehn, Larry see Gupton, Vol. 6,

98. Jain, Kamal, Vol. 10, 46. Kennedy, Robert E. see Cooper,

Vol. 6, 160. Kennedy, Robert E. see Cooper,

Vol. 7, 40. Kennedy, Robert E. see Cooper,

Vol. 7, 87. Kennedy, Robert E. and Curtis

Cooper, Vol. 7, 139. Kennedy, Robert E. see Cooper,

Vol. 7, 140. Kennedy, Robert E. see Cooper,

Vol. 8, 90.

Kennedy, Robert E. see Rabi-

nowitz, Vol. 8, 90. Palmer, Leonard L., Vol. 6, 47. Palmer, Leonard L., Vol. 6, 160. Palmer, Leonard L., Vol. 8, 136. Palmer, Leonard L., Vol. 9, 34. Palmer, Leonard L., Vol. 9, 106. Rabinowitz, Stanley, Vol. 6, 47. Rabinowitz, Stanley, Vol. 8, 36. Rabinowitz, Stanley, Robert E.

Kennedy and Curtis Cooper, Vol. 8, 90.

Rapoport, Alan H., Vol. 7, 139. Rapoport, Alan H., Vol. 8, 89. Ridenhour, Jim see Gupton, Vol. 6,

98. Risteski, Ice B., Vol. 10, 176. Risteski, Ice B., Vol. 10, 176. Sadek, Jawad see Euler, Russell,

Vol. 10, 47. Sadek, Jawad see Euler, Russell,

Vol. 10, 118. Skow, Donald P., Vol. 7, 88. Sriskandarajah, J., Vol. 6, 98. Sriskandarajah, J., Vol. 7, 87. Taylor, James H., Vol. 7, 140. Timm, Mathew, Vol. 9, 185. Wheatley, W. F. and James

Etheridge, Vol. 7, 88.

Problems Solved Ahuja, Mangho, Vol. 9, 196 Ahuja, Mangho see Woods, Vol. 6,

167. Azarian, Mohammad K., Vol. 6,

51. Azarian, Mohammad K., Vol. 6,

161. Azarian, Mohammad K., Vol. 10,

128. Azarian, Mohammad K., Vol. 10,

183. Azarian, Mohammad K. see

Chance, Vol. 6, 99. Bang, Seung-Jin see Doucette, Vol.

6, 49. Bang, Seung-Jin, Curtis Cooper,

and Robert E. Kennedy, Vol. 6, 54.

Bang, Seung-Jin see Chance, Vol.

6, 99. Belshoff, Richard see Reid, Les.

Vol. 9, 186. Bergum, Gerald E., Vol. 7, 47. Bergum, Gerald E., Vol. 9, 114. Bruckman, Paul S., Vol. 10, 178. Bruening, James T., Vol. 8, 157. Bruening, James T., Vol. 9, 121. Bruening, James T., Vol. 9, 189. Bruening, James T., Vol. 9, 193. Bruening, James T., Vol. 9, 195. Bruening, James T., see Prielipp,

Vol. 9, 111. Bruening, James T. and Kan-

dasamy Muthuvel, Vol. 10, 129. Bruton, Gregory see Prielipp, Vol.

6, 49. Chance, Joseph E. see Doucette,

Vol. 6, 49. Chance, Joseph E., Vol. 6, 50. Chance, Joseph E., Vol. 6, 53. Chance, Joseph E., Vol. 10, 181.

Chance, Joseph E., Seung-Jin Bang, and Mohammad K. Azarian, Vol. 6, 99.

Chance, Joseph E. see Tinsley, Vol.

6, 109. Chance, Joseph E. see Woods, Vol.

6, 167. Chance, Joseph E. see Prielipp,

Vol. 7, 43. Cooper, Curtis and Robert E.

Kennedy, Vol. 9, 112. Cooper, Curtis see Bang, Vol. 6,

54. Cooper, Curtis see Dence, Vol. 8,

93. Cooper, Curtis see Kennedy, Vol.

8, 153. Davenport, Kenneth see Prielipp,

Vol. 9, 111. Davenport, Kenneth see Dence,

Joseph B., Vol. 10, 119. Davenport, Kenneth see Euler,

Russell, Vol. 10, 126. Dawson, Bryan, Vol. 9, 197. Dence, Joseph B., Vol. 7, 141. Dence, Joseph B., Vol. 7, 144. Dence, Joseph B., Vol. 7, 145. Dence, Joseph B., Vol. 8, 137. Dence, Joseph B., Vol. 8, 163. Dence, Joseph B., Vol. 8, 167. Dence, Joseph B., Vol. 9, 37. Dence, Joseph B., Vol. 9, 45. Dence, Joseph B., Vol. 9, 58. Dence, Joseph B., Vol. 9, 62. Dence, Joseph B., Vol. 10, 131. Dence, Joseph B. see Woods, Vol.

6, 167. Dence, Joseph B. see Prielipp, Vol.

7, 143. Dence, Joseph B. see Prielipp, Vol.

7, 146.

Dence, Joseph B., Curtis Cooper and Robert E. Kennedy, Vol. 8, 93.

Dence, Joseph B. see Prielipp, Vol.

8, 95. Dence, Joseph B. see Euler, Vol. 8,

96. Dence, Joseph B. see Prielipp, Vol.

9, 111. Dence, Joseph B. see Prielipp, Vol.

10, 49. Dence, Joseph B. see Euler, Russel-

l, Vol. 10, 60. Dence, Joseph B., Joe Howard,

Kandasamy Muthuvel, and Kenneth Davenport, Vol. 10, 119. Dence, Joseph B. and Carl Libis,

Vol. 10, 124. Doucette, Robert L., Donald P. Skow, Kandasamy Muthuvel, Joseph E. Chance, Seung-Jin Bang, Joe Howard, J. Sriskandarajah, and Russell Euler, Vol. 6, 49.

Euler, Russell, Vol. 8, 42. Euler, Russell, Vol. 9, 48. Euler, Russell, Vol. 9, 56. Euler, Russell, Vol. 9, 111. Euler, Russell see Woods, Vol. 6,

167. Euler, Russell see Freitag, Vol. 6,

167. Euler, Russell see Prielipp, Vol. 7,

43. Euler, Russell see Prielipp, Vol. 7,

143. Euler, Russell see Prielipp, Vol. 7,

146. Euler, Russell see Prielipp, Vol. 8,

95. Euler, Russell, Joseph B. Dence, Lawrence Somer, J. Sriskandarajah, Herta T. Freitag and Donald P. Skow, Vol. 8, 96.

Euler, Russell and Jawad Sadek,

Vol. 8, 165. Euler, Russell see Prielipp, Vol. 9,

111. Euler, Russell, Jawad Sadek, and

Joseph B. Dence, Vol. 10, 60. Euler, Russell, Jawad Sadek, and

Kenneth Davenport, Vol. 10, 126.

Euler, Russell and Jawad Sadek,

Vol. 10, 179. Flanigan, Frank J. see Kuenzi, Vol.

6, 100. Flanigan, Frank J., Vol. 6, 102. Flanigan, Frank J., Vol. 8, 164. Flanigan, Frank J., Vol. 9, 38. Freitag, Herta T. and Russell Eu-

ler, Vol. 6, 167. Freitag, Herta T. see Prielipp, Vol.

7, 143. Freitag, Herta T. see Prielipp, Vol.

7, 147. Freitag, Herta T., Vol. 8, 39. Freitag, Herta T., Vol. 9, 51. Freitag, Herta T., Vol. 9, 191. Freitag, Herta T. see Prielipp, Vol.

8, 95. Freitag, Herta T. see Euler, Vol. 8,

96. Freitag, Herta T. see Howard, Vol.

8, 96. Hall, Leon see Kuenzi, Vol. 6, 104. Hall, Leon see Woods, Vol. 6, 167. Howard, Joe, Vol. 10, 127. Howard, Joe, Vol. 10, 183. Howard, Joe see Doucette, Vol. 6,

49. Howard, Joe see Prielipp, Vol. 7,

43.

Howard, Joe see Prielipp, Vol. 7,

146. Howard, Joe see Prielipp, Vol. 7,

147. Howard, Joe see Prielipp, Vol. 8,

95. Howard, Joe and Herta T. Freitag,

Vol. 8, 96. Howard, Joe see Prielipp, Vol. 10,

51. Howard, Joe, and Joseph Wiener,

Vol. 10, 59. Howard, Joe see Dence, Joseph B.,

Vol. 10, 119. Hoehn, Larry see Prielipp, Vol. 7,

43. Johnson, Rob, Vol. 8, 42. Johnson, Rob and Lawrence Somer,

Vol. 8, 43. Kennedy, Robert E. see Bang, Vol.

6, 54. Kennedy, Robert E. see Dence, Vol.

8, 93. Kennedy, Robert E. see Cooper,

Vol. 8, 152. Kennedy, Robert E. see Cooper,

Vol. 9, 112. Koker, John see Prielipp, Vol. 7,

43. Kuenzi, N. J., Vol. 6, 165. Kuenzi, N. J., Vol. 7, 42. Kuenzi, N. J., Vol. 8, 44. Kuenzi, N. J., Vol. 8, 98. Kuenzi, N. J., Vol. 8, 100. Kuenzi, N. J., Vol. 8, 101. Kuenzi, N. J., Vol. 9, 107. Kuenzi, N. J., Vol. 9, 193. Kuenzi, N. J. and Frank J. Flani-

gan, Vol. 6, 100. Kuenzi, N. J., Leon Hall, and

Rhonda McKee, Vol. 6, 104.

Kuenzi, N. J. and Bob Prielipp,

Vol. 6, 163. Kuenzi, N. J. see Prielipp, Vol. 8,

43. Kuenzi, N. J., John Oman, and

Bob Prielipp, Vol. 10, 56. Kuznetsova, Elizaveta see Prielipp,

Vol. 7, 146. Leong, Thomas C., Vol. 9, 36. Leong, Thomas C., Vol. 9, 42. Leong, Thomas C., Vol. 9, 115. Leong, Thomas C. see Prielipp,

Vol. 9, 111. Libis, Carl, Vol. 10, 120. Libis, Carl see Prielipp, Vol. 10,

182. Masuda, Jerry see Prielipp, Vol.

10, 182. McKee, Rhonda see Kuenzi, Vol. 6,

104. Muthuvel, Kandasamy, Vol. 6, 101. Muthuvel, Kandasamy, Vol. 8, 38. Muthuvel, Kandasamy, Vol. 9, 39. Muthuvel, Kandasamy see

Doucette, Vol. 6, 49. Muthuvel, Kandasamy see Prielipp,

Vol. 6, 163. Muthuvel, Kandasamy see Dence,

Joseph B., Vol. 10, 119. Muthuvel, Kandasamy see Bruen-

ing, James T., Vol. 10, 129. Nanyes, Ollie, Vol. 6, 102. Oman, John see Prielipp, Vol. 10,

51. Oman, John see Kuenzi, N. J., Vol.

10, 56. Palmer, Leonard L., Vol. 10, 130. Palmer, Leonard L. see Prielipp,

Vol. 7, 43. Palmer, Leonard L. see Prielipp,

Vol. 7, 146.

Palmer, Leonard L. see Prielipp,

Vol. 10, 49. Palmer, Sherri see Prielipp, Vol. 6,

49. Prielipp, Bob, Vol. 7, 44. Prielipp, Bob, Vol. 7, 142. Prielipp, Bob, Vol. 8, 45. Prielipp, Bob, Vol. 8, 45. Prielipp, Bob, Vol. 8, 92. Prielipp, Bob, Vol. 8, 160. Prielipp, Bob, Vol. 9, 194. Prielipp, Bob, Vol. 10, 54. Prielipp, Bob, Vol. 10, 58. Prielipp, Bob, Vol. 10, 121. Prielipp, Bob, Vol. 10, 125. Prielipp, Bob, Gregory Bruton, and Sherri Palmer, Vol. 6, 49. Prielipp, Bob see Kuenzi, Vol. 6,

163. Prielipp, Bob, Kandasamy Muthuvel, J. Sriskandarajah, Vol. 6, 166. Prielipp, Bob, Larry Hoehn, John Koker, Joe Howard, Lawrence Somer, Joseph E. Chance, Russell Euler, Donald P. Skow, and Leonard L. Palmer, Vol. 7, 43.

Prielipp, Bob, Russell Euler, Lawrence Somer, Joseph B. Dence, and Herta T. Freitag, Vol. 7, 143.

Prielipp, Bob, Joseph B. Dence, Bella Wiener, Russell Euler, Joe Howard, Elizaveta Kuznetsova, Lawrence Somer, R. P. Sealy, and Leonard L. Palmer, Vol. 7, 146. Prielipp, Bob see Wiener, Vol. 7,

147. Prielipp, Bob and N. J. Kuenzi,

Vol. 8, 43. Prielipp, Bob, Russell Euler, Joe Howard, Joseph B. Dence, Lawrence Somer, Gayla Singleton, J. Sriskandarajah, Herta T. Freitag and Donald P. Skow, Vol. 8, 95.

Prielipp, Bob and Herta T. Freitag,

Vol. 8, 96. Prielipp, Bob, Thomas C. Leong,

James T. Bruening, Joseph B. Dence, Russell Euler, Jawad Sadek, Lawrence Somer, and Kenneth Davenport, Vol. 9, 111.

Prielipp, Bob, Joseph B. Dence, and Leonard L. Palmer, Vol. 10, 49.

Prielipp, Bob, John Oman, Joe Howard, and Donald P. Skow, Vol. 10, 51.

Prielipp, Bob see Kuenzi, N. J.,

Vol. 10, 56. Prielipp, Bob, Carl Libis, Jerry Ma-

suda, Ice B. Risteski, Vol. 10, 182.

Rabinowitz, Stanley, Vol. 6, 52. Radhakrishnan, Srikant, Vol. 8,

139. Rapoport, Alan H., Vol. 8, 102. Rapoport, Alan H., Vol. 8, 153. Rapoport, Alan H., Vol. 9, 110. Redmond, Don, Vol. 9, 113. Reid, Les and Richard Belshoff,

Vol. 9, 186. Risteski, Ice B., Vol. 10, 184. Risteski, Ice B. see Prielipp, Vol.

10, 182. Sadek, Jawad see Euler, Vol. 8,

165. Sadek, Jawad see Prielipp, Vol. 9,

111. Sadek, Jawad see Euler, Russell,

Vol. 10, 60. Sadek, Jawad see Euler, Russell,

Vol. 10, 126.

Sadek, Jawad see Euler, Russell,

Vol. 10, 179. Sealy, R. P. see Prielipp, Vol. 7,

146. Singleton, Gayla see Prielipp, Vol.

8, 95. Skow, Donald P., Vol. 6, 50. Skow, Donald P., Vol. 9, 194 Skow, Donald P. see Doucette, Vol.

6, 49. Skow, Donald P. see Prielipp, Vol.

7, 43. Skow, Donald P. see Prielipp, Vol.

8, 95. Skow, Donald P. see Euler, Vol. 8,

96. Skow, Donald P. see Prielipp, Vol.

10, 51. Smith, Ronald K., Vol. 8, 97. Somer, Lawrence, Vol. 9, 37. Somer, Lawrence see Prielipp, Vol.

7, 43. Somer, Lawrence see Prielipp, Vol.

7, 143. Somer, Lawrence see Prielipp, Vol.

7, 146. Somer, Lawrence see Johnson, Vol.

8, 43. Somer, Lawrence see Prielipp, Vol.

8, 95. Somer, Lawrence see Euler, Vol. 8,

96. Somer, Lawrence see Prielipp, Vol.

9, 111. Sriskandarajah, J., Vol. 6, 161. Sriskandarajah, J., Vol. 6, 166. Sriskandarajah, J. see Doucette,

Vol. 6, 49. Sriskandarajah, J. see Tinsley, Vol.

6, 109. Sriskandarajah, J. see Prielipp, Vol.

8, 95.

Sriskandarajah, J. see Euler, Vol.

8, 96. Timm, Mathew, Vol. 10, 185. Tinsley, Alvin, Joseph E. Chance,

and J. Sriskandarajah, Vol. 6, 109.

Widmer, Lamarr, Vol. 8, 38. Widmer, Lamarr, Vol. 9, 60. Wiener, Joseph, Vol. 10, 48. Wiener, Joseph see Howard, Joe,

Vol. 10, 59. Wiener, Bella see Prielipp, Vol. 7,

146. Wiener, Bella, Bob Prielipp, Joe

Howard, and Herta T. Freitag, Vol. 7, 147.

Woods, Dale, Vol. 8, 43. Woods, Dale, Vol. 9, 193. Woods, Dale, Vol. 9, 193. Woods, Dale, Leon Hall, Joseph E.

Chance, Mangho Ahuja, Joseph B. Dence, and Russell Euler, Vol. 6, 167. æ